Thursday July 7,1994
Work commenced with a general cleaning of the H/20-22 area that was being excavated yesterday. Scraping with trowels revealed no difference in the soil of H/20 ane the soil of H/21. In other words it is not yet possible to discern the western edge of T-6b. Once the cleaning was accomplished, pick and shovel were used to continue excavating this area.
As H/20-22 was excavated, it appeared that less pottery and tile were bing recovered at 30+cm than yesterday at 10-30cm. This relative paucity of material continued up to a depth of 45cm, where work stopped for the day.
tile= less than 1/8 casetta
pottery= 1/4 liter box
|Property or Relation||Value(s)|
[Standard: Dublin Core Terms]
Open Context editors work with data contributors to annotate datasets to shared vocabularies, ontologies, and other standards using 'Linked Open Data' (LOD) methods.
The annotations presented above approximate some of the meaning in this contributed data record to concepts defined in shared standards. These annotations are provided to help make datasets easier to understand and use with other datasets.
To the extent to which copyright applies, this content
carries the above license. Follow the link to understand specific permissions
Required Attribution: Citation and reference of URIs (hyperlinks)